5. A shoes manufacturing company finds from the sales data of the other manufactures of similar product that the demand functions for the shoes can be expressed as
Q=10,000-20 P
Find:-
a. Number of shoes sold when P=$300
b. Price for selling 5400 pair of shoes
c. Price of zero sales
d. Point elasticity of demand at price $500
1
Expert's answer
2020-12-10T14:35:50-0500

a)     Q=10000-20p

Q=10,000-20(300)

Q=4,000 shoes

b)    5,400=10,000-20p

20p=10,000-5,400

P=$230

c)   0=10,000-20p

20p=10,000

P=$500

d)  Ep=(∆Q/∆P) ×P/Q

∆Q/∆P=-20/1=-20

Q=10,000-20×500

Q=0

Ep=-20(500/0)

Ep=∞

At price $500 demand is perfectly elastic